The following table shows the number of U.S. dollars required to buy one Mexican peso and the number of U.S. dollars required to buy one Japanese yen between June 1, 2016, and September 1, 2016. Use this table to answer the following questions. Date June 1, 2016 July 1, 2016 August 1, 2016 September 1, 2016 U.S. Dollars Required to Buy 1 Mexican Peso 0.0536 0.0537 0.0541 0.0520 U.S. Dollars Required to Buy 1 Japanese Yen 0.0095 0.0096 0.0099 0.0098 On July 1, 2016, Brandy paid 450 Mexican pesos for a manicure in Playa del Carmen, Mexico. Because she only had Japanese yen, she needed to exchange her Japanese yen at a nearby money- exchange kiosk in order to get the 450 Mexican pesos she used to pay the manicurist. The kiosk is able to buy and sell U.S. dollars, Mexican pesos, and Japanese yen at the exchange rates shown in the table. Approximately how much did the manicure cost Brandy in terms of Japanese yen? O 2.921 0.3386 0.3243 O 2.517 472,532
Answer #1


On July 1, 2016, exchange rate between US dollars and Mexican Peso is as follows -

1 Mexican Peso = $0.0537

So,

450 Mexican Peso = $0.0537 * 450 = $24.165

On July 1, 2016, exchange rate between US dollars and Japanese Yen is as follows -

1 Japanese Yen = $0.0096

$1 = (1/0.0096) Japanese Yen

$24.165 = (1/0.0096) * 24.165 = 2,517.18 Japanese Yen

So,

The manicure would cost Brandy 2,517 Japanese Yen.

Hence, the correct answer is the option (D) [2,517]